Consider the market for pop. Use the demand and supply equations.

Qd

​=

100minus−20P

Qs

​= 5P

What is the marginal benefit of the

10th

can of​ pop?
Answer #1

Marginal Benefit is same as the Inverse Demand curve.

So, Marginal Benefit = Inverse Demand curve

As, Demand curve :-  

Q​​​​​​d​​​​​ = 100 - 20P

​​​​​​20P = 100 - Q​​​​​​d

P = (100 - Q​​​​​d ) / 20

Inverse Demand curve :- P = (100 - Q​​​​​d ) / 20

So, At Q = 10

P = (100 - 10) / 20

P = 90/20 = 4.5

​​​​​​So, Marginal Benefit is 4.5 from the 10th can of pop.

​​​​​